The width of a rectangle is 5 feet more than its length. The area of the rectangle is 36 sqaure feet. What is the length, in feet, of this rectangle?
Answer by richwmiller(17219)

W=L+5
W+L=36
l*(l+5)=36
l^2+5l=36
L^2+5L-36=0
(l+9)(l-4)=0
l=4 w=9